Marine fisheries is an established company which is looking to expand its fishing interests by purchasing a 100%interest in shark bait.The management of marine fisheries believes that the expected return from the acquisition of shark bait are dependent on the state of the economy. State of the economy.favourable, neutral, unfavorable. Probability of occurrence 0.3,0.4,0.3 .marine fisheries 16%,10%,2%.shark bait 20%,12%,0%.The market 14%, 8%,6%. Book value in millions R12m,R8m and R0. Market value in millions R8m,R12m and R0 Standard deviation of return 5,4%,7,8%and3.2%. Covariance with the market 0.0024 0,0023 and 0.the risk free rate is 15%and the is no company or personal taxation. Required:determine whether marine fisheries should acquire shark bait using the portfolio management theory.
1
Expert's answer
2020-08-03T16:05:15-0400

"r1=0.15+0.0024*(0.15-0.14)=0.150024"

"r2=0.15+0.0023*(0.15-0.08)=0.150161"

"r3=0.15"

r1-profitability (favourable); r2-profitability (neutral); r3-profitability (unfavorable)

"Ra1=0.3*(0.16+0.2)+0.4*(0.1+0.12)+0.3*0.02=0.202=20.2\%"

"Ra2=0.3*(0.16+0.150024)+0.4*(0.1+0.150161)+0.3*(0.02+0.15)=0.2440716=24.40716\%"

Ra1- Risk-adjusted income if the shark bait will be purchased

Ra2- Risk-adjusted income if the shark bait will not be purchased and Marine fisheries will invest in financial assets

Ra1<Ra2

the buying of shark bait is not profitable